# ToMoE: Converting Dense Large Language Models to Mixture-of-Experts through Dynamic Structural Pruning

### v1

@@ 4.4 Analysis of ToMoE
Training Dynamics. In Fig.3, we visualize the training dynamics under different values of $p$. Across all $p$, the knowledge distillation loss $\mathcal{L}$ (Fig.3(a)), the parameter regularization loss $\mathcal{R}_{\text{P}}$, and the union of experts regularization loss $\mathcal{R}_{\text{U}}$ decrease over the course of training. Notably, the parameter regularization loss quickly drops to 00 in the early stages of training, while using a smaller $p$ requires more iterations. The peak of the union of experts regularization loss increases when using smaller values of $p$, indicating that the initial solution tends to only cover a small portion of the dense model. Regarding the load balancing loss, it oscillates around 0.15, demonstrating that ToMoE maintains a relatively balanced load distribution during the training process.
Ablation Study. In Fig. 6, we present the average zero-shot task performance under different settings. For $p=0.4$ and $p=0.5$, replacing the knowledge distillation loss with the language modeling loss significantly impacts performance. At $p=0.4$, removing $\mathcal{R}_{\text{U}}$ also results in a substantial performance drop, whereas the impact is much smaller at $p=0.5$. We hypothesize that this difference arises because reducing $p$ makes the learning process more challenging. Without the guidance provided by $\mathcal{R}_{\text{U}}$, the model struggles to effectively utilize the parameters of the original model. Additionally, the choice of dataset affects performance, particularly when switching to the WikiText dataset. This demonstrates that a mixing dataset is beneficial to the overall performance.
Other Analysis. (1). We present the width of our ToMoE model for Qwen-2.5 7B in Fig. 4. We can see that the layer-wise configuration is highly non-uniform, showing that our method can flexibly set the width of different layers and operations. (2). In Fig. 5(a), we show that the union of experts is close to the full model capacity even though the width of experts across different layers is highly non-uniform, demonstrating the effectiveness of our loss design. (3). In Fig. 5(b), we show the costs of different learning-based methods in terms of US dollars. ToMoE has a similar cost compared to DISP-LLM with LLaMA-2 7B and 13B models, and both of them are much cheaper than LLM Surgeon. (4). Finally, we visualize the expert selection for LLaMA-2 7B in Tab. 4. We can see that each expert aligns syntax rather than semantic meanings resembling the observations in (Jiang et al., 2024).
@@ 5 Conclusion
Impact Statement
This paper introduces a method to reduce the active parameters of LLMs during inference, thereby lowering their computational costs. By improving efficiency, it can lead to a decrease in carbon emissions, contributing to environmental sustainability and benefiting society and industry alike. Additionally, reducing the costs associated with LLMs could make their services more affordable, broadening access to powerful AI tools and ensuring that AI development remains inclusive and widely accessible.
@@ A.1 Module Configurations
HyperNetwork Input $z$rightarrow Bi-GRU(32,64)
@@
We present the details of trainable modules in Tab. 5. In short, we project the input tokens to a low-dimensional space and add them to the output of the HyperNetwork. The input $z$ to the HyperNetwork are fixed random vectors of size $N\times 32$ sampled from Normal Distribution. Except for the HyperNetwork, other individual trainable modules are created for each MHA and MLP layer. If we have $L$ blocks, then we will have $L$ $\text{Proj}_{\text{E}}^{\text{\tiny MHA}}$, $L$ $\text{Proj}_{\text{D}}^{\text{\tiny MHA}}$ with output size of $\frac{d}{H}$, $L$ $\text{Proj}_{\text{D}}^{\text{\tiny MHA}}$ with output size of $\frac{d}{2H}$ $L$, $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}$ and $L$ Router layers. Notations of $d$, $H$, $d_{\text{mid}}$, and $N$ are already defined in Sec. 2.
@@
${}\mathbf{s}_{\scriptstyle{\mathbf{e}}}=\text{ST-GS}(\text{Proj}_{\text{D}}^{\text{\tiny MLP}}(\mathbf{E})),$ (11)
where ST-GS again is the Straight-Through Gumbel-Sigmoid function. $\mathbf{s}_{\scriptstyle{\mathbf{e}}}\in\{0,1\}^{N\times d_{\text{mid}}}$ is the resulting binary vectors to select experts from the dense model. Once $\mathbf{s}_{\scriptstyle{\mathbf{e}}}$ is generated, it can be reused and thus we no longer need $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}$. Let $\mathbf{S}_{\scriptstyle{\mathbf{e}}}^{i}=\text{Diag}(\mathbf{s}_{\scriptstyle{\mathbf{e}}}^{i}),\ i=1,\cdots,N$. Similarly, we use $\hat{\mathbf{s}}_{\scriptstyle{\mathbf{e}}}^{i}\in\Re^{d_{\text{mid}}\times d_{\text{mid}}^{\prime}}$ to represent the actual column or row selection matrix by removing zero columns or rows, where $d_{\text{mid}}^{\prime}<d_{\text{mid}}$ and it is the width of each expert. The $i$th expert can be represented as:
$f_{\text{MLP}}^{i}(\mathbf{X})=\sigma(\mathbf{X}\mathbf{W}_{G}\hat{\mathbf{S}}_{\scriptstyle{\mathbf{e}}}^{i})\odot(\mathbf{X}\mathbf{W}_{U}\hat{\mathbf{S}}_{\scriptstyle{\mathbf{e}}}^{i})\hat{\mathbf{S}}_{\scriptstyle{\mathbf{e}}}^{i}\mathbf{W}_{D}.\$ (12)
After ToMoE, given the result of the routing function $\mathbf{G}={\text{ST-G}}(\text{Router}(\mathbf{X}))$, the MLP calculation with MoE can be written as:
@@
Another question is how many parameters we need after introducing Top-K routing for MHA layers and Top-1 routing for MLP layers. Analytically, the additional parameters can be calculated by $L\times d\times 128+L\times 128\times(\frac{d}{H})+L\times d\times N$. Let’s use LLaMA-2 7B as an example, $L=32$, $d=4096$, $N=8$, $H=32$, the additional parameters are $1\times 32\times 4096\times 128+32\times 128\times(128)+32\times 4096\times 8=0.0184$B. This is equivalent to $0.27\%$ of the total parameters of the LLaMA-2 7B model, and thus the additional parameter is not significant compared to the original number of parameters.
@@
Rotary Position Embedding (RoPE) (Su et al., 2024) is a popular positional encoding method and it is regularly used in LLMs like LLaMA (Touvron et al., 2023b). RoPE divides the $\frac{d}{H}$ dimensional space into $\frac{d}{2H}$ sub-spaces, and they are applied on query and key. This means that if we want to perform head dimension pruning for query and key matrices, we need to follow the sub-spaces resulting from RoPE and make these two sub-spaces share the same pruning mask $\mathbf{s}_{0}$, and the final pruning mask for query and key is $\mathbf{s}_{0}^{\prime}=[{\mathbf{s}_{0}}_{[1:\frac{d}{2H}]},{\mathbf{s}_{0}}_{[1:\frac{d}{2H}]}]$, and clearly the size of ${\mathbf{s}_{0}}_{[1:\frac{d}{2H}]}$ is $\frac{d}{2H}$. In short, we simply select the first half of elements from $\mathbf{s}_{0}$ and repeat it twice to have the final pruning decision. We also found that applying dynamic pruning for query and key matrices along the head dimension is difficult and unreasonable since different tokens may have different positions after pruning. It becomes a problem when calculating the inner product between the query and key matrices given different tokens.
By applying head dimension pruning, our method also does not need to be modified when facing different attention mechanisms like GQA (Grouped-Query Attention) (Ainslie et al., 2023) and MQA (Multi-Query Attention) (Shazeer, 2019).
@@
70% SliceGPT (Ashkboos et al., 2024) 60.27 36.18 69.42 64.09 49.74 55.94
LLM Surgeon (van der Ouderaa et al., 2024) 69.74 40.27 76.50 68.67 71.52 65.34
DISP-LLM (Gao et al., 2024) 63.80 39.42 74.43 66.85 70.86 63.07
DISP-LLM Alpaca (Gao et al., 2024) 68.98 44.28 77.31 67.32 68.98 65.59
MoDeGPT-Alpaca (Lin et al., 2024) 70.24 41.47 77.15 71.27 71.84 66.39
@@
60% SliceGPT (Ashkboos et al., 2024) 48.99 32.51 63.17 56.75 39.85 48.25
LLM Surgeon (van der Ouderaa et al., 2024) 63.80 37.12 73.16 65.75 65.64 60.94
DISP-LLM (Gao et al., 2024) 62.67 35.63 73.39 62.67 65.86 60.04
DISP-LLM Alpaca (Gao et al., 2024) 66.79 42.75 75.30 64.25 67.52 63.32
MoDeGPT-Alpaca (Lin et al., 2024) 63.72 38.82 71.87 66.30 62.10 60.56
@@
50% LLM Surgeon (van der Ouderaa et al., 2024) 56.19 37.12 68.87 63.22 56.19 56.32
DISP-LLM (Gao et al., 2024) 58.27 36.87 68.67 59.27 57.18 54.50
DISP-LLM Alpaca (Gao et al., 2024) 55.72 37.54 72.20 59.59 62.39 57.49
@@ A.3 Details of Gumbel-Softmax and Gumbel-Sigmoid
The Gumbel-Softmax function (Jang et al., 2016) allows for differentiable sampling from a categorical distribution. Given logits $\mathbf{x}$, the Gumbel-Softmax sample $\mathbf{y}$ is computed as:
@@
where each element of $\mathbf{g}$ is drawn from $\text{Gumbel}(0,1)$, and $\tau$ is the temperature parameter that controls the smoothness of the distribution. Combining Gumbel-Softmax with the Straight-Through gradient Estimator (Bengio et al., 2013), we have the following equation:
[…]
where $D=\{1,2,\cdots,N\}$, $N$ again is the number of experts in our setting, and one-hot will assign $1$ corresponding to the position of the maximum value in $\frac{\mathbf{x}+\mathbf{g}}{\tau}$ and assign 00 to other positions.
The Gumbel-Sigmoid function is a special case of the Gumbel-Softmax function, designed for binary distributions. Given logits $\mathbf{x}$, the Gumbel-Sigmoid sample $y$ is computed as:
$y=\text{sigmoid}\left(\frac{\mathbf{x}+\mathbf{g}}{\tau}\right),$
@@
[…]
where $b$ is a constant bias in our implementation and it ensures that all experts start from the whole model, $\text{round}(\cdot)$ will round the input values to the nearest integer, and in our case, it rounds inputs to 00 or $1$. For all experiments, we set $b=3.0$ in Eq. 15, and we set $\tau=0.4$ for Eq. 14 and Eq. 15.
@@
DISP-LLM (Gao et al., 2024) 68.18 44.11 74.86 67.09 62.93 63.43
75% SliceGPT (Ashkboos et al., 2024) 53.70 31.66 69.21 65.35 52.40 54.46
@@
DISP-LLM (Gao et al., 2024) 65.93 43.34 74.27 65.11 59.95 61.72
70% SliceGPT (Ashkboos et al., 2024) 53.03 30.29 65.94 63.14 47.56 51.99
@@
Appendix B More Details of the Loss Design
B.1 Implementation of the Self-Knowledge Distillation
During the ToMoE learning process, we freeze the parameters of the original model. This approach offers the additional benefit of enabling self-knowledge distillation without the need to load an extra model. In Lst. LABEL:lst:self-kd, we present the pseudo-code for the self-knowledge distillation process. In summary, we first disable the trainable modules associated with ToMoE and compute the output logits from the original model. Next, we re-enable the trainable modules for ToMoE and perform a regular forward pass. The logits from the original model are then used to guide the learning of ToMoE.

### v2

@@ 4.4 Analysis of ToMoE
Training Dynamics. In Fig.3, we visualize the training dynamics under different values of $p$. Across all $p$, the knowledge distillation loss $\mathcal{L}$ (Fig.3(a)), the parameter regularization loss $\mathcal{R}_{\text{P}}$, and the union of experts regularization loss $\mathcal{R}_{\text{U}}$ decrease over the course of training. Notably, the parameter regularization loss quickly drops to 0 in the early stages of training, while using a smaller $p$ requires more iterations. The peak of the union of experts regularization loss increases when using smaller values of $p$, indicating that the initial solution tends to only cover a small portion of the dense model. Regarding the load balancing loss, it oscillates around 0.15, demonstrating that ToMoE maintains a relatively balanced load distribution during the training process.
Ablation Study. In Fig. 6(c), we present the average zero-shot task performance under different settings. For $p=0.4$ and $p=0.5$, replacing the knowledge distillation loss with the language modeling loss significantly impacts performance. At $p=0.4$, removing $\mathcal{R}_{\text{U}}$ also results in a substantial performance drop, whereas the impact is much smaller at $p=0.5$. We hypothesize that this difference arises because reducing $p$ makes the learning process more challenging. Without the guidance provided by $\mathcal{R}_{\text{U}}$, the model struggles to effectively utilize the parameters of the original model. Additionally, the choice of dataset affects performance, particularly when switching to the WikiText dataset. This demonstrates that a mixing dataset is beneficial to the overall performance.
Other Analysis. (1). Fig. 5 presents the width of our ToMoE model for Qwen-2.5 7B, which shows the layer-wise configuration is highly non-uniform. It demonstrates that our method can flexibly set the width of different layers and operations. (2). Fig. 6(a) shows that the union of experts is close to the full model capacity, even though the width of experts across different layers is highly non-uniform, demonstrating the effectiveness of our loss design. (3). Fig. 6(b) plots the costs of different learning-based methods in terms of US dollars. ToMoE costs similarly compared to DISP-LLM with LLaMA-2 7B and 13B models, and both of them are much cheaper than LLM Surgeon. (4). Fig. 4 shows the token allocation across experts on the Wikitext dataset. We observe that the early and late layers exhibit relatively balanced expert utilization, while the middle layers have certain experts activated more frequently. (5). Finally, we visualize the expert selection for LLaMA-2 7B in Tab. 7. We can observe that each expert aligns syntax rather than semantic meanings, resembling the observations in (jiang2024mixtral).
@@ A.1 Module Configurations
We present the details of trainable modules in Tab. 8. In short, we project the input tokens to a low-dimensional space and add them to the output of the HyperNetwork. The inputs $z$ to the HyperNetwork are fixed random vectors of size $N\times 32$ sampled from a Normal Distribution. Except for the HyperNetwork, other individual trainable modules are created for each MHA and MLP layer. If we have $L$ blocks, then we will have $L$ $\text{Proj}_{\text{E}}^{\text{\tiny MHA}}$, $L$ $\text{Proj}_{\text{D}}^{\text{\tiny MHA}}$ with output size of $\frac{d}{H}$, $L$ $\text{Proj}_{\text{D}}^{\text{\tiny MHA}}$ with output size of $\frac{d}{2H}$ $L$, $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}$ and $L$ Router layers. Notations of $d$, $H$, $d_{\text{mid}}$, and $N$ are already defined in Sec. 2.
@@
HyperNetwork Input $z$ $\rightarrow$ Bi-GRU(32,64)
@@
$\mathbf{s}_{\scriptstyle{\mathbf{e}}}=\text{ST-GSig}(\text{Proj}_{\text{D}}^{\text{\tiny MLP}}(\mathbf{E})),$ (11)
where ST-GSig again is the Straight-Through Gumbel-Sigmoid function. $\mathbf{s}_{\scriptstyle{\mathbf{e}}}\in\{0,1\}^{N\times d_{\text{mid}}}$ is the resulting binary vectors to select experts from the dense model. Once $\mathbf{s}_{\scriptstyle{\mathbf{e}}}$ is generated, it can be reused, and thus we no longer need $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}$. Let $\mathbf{S}_{\scriptstyle{\mathbf{e}}}^{i}=\text{Diag}(\mathbf{s}_{\scriptstyle{\mathbf{e}}}^{i}),\ i=1,\cdots,N$. Similarly, we use $\hat{\mathbf{s}}_{\scriptstyle{\mathbf{e}}}^{i}\in\Re^{d_{\text{mid}}\times d_{\text{mid}}^{\prime}}$ to represent the actual column or row selection matrix by removing zero columns or rows, where $d_{\text{mid}}^{\prime}<d_{\text{mid}}$ and it is the width of each expert. The $i$th expert can be represented as:
$f_{\text{MLP}}^{i}(\mathbf{X})=\sigma(\mathbf{X}\mathbf{W}_{G}\hat{\mathbf{S}}_{\scriptstyle{\mathbf{e}}}^{i})\odot(\mathbf{X}\mathbf{W}_{U}\hat{\mathbf{S}}_{\scriptstyle{\mathbf{e}}}^{i})\hat{\mathbf{S}}_{\scriptstyle{\mathbf{e}}}^{i}\mathbf{W}_{D}.\\$ (12)
After ToMoE, given the result of the routing function $\mathbf{G}={\text{ST-GSmax}}(\text{Router}(\mathbf{X}))$, the MLP calculation with MoE can be written as:
@@
Another question is how many parameters we need after introducing Top-K routing for MHA layers and Top-1 routing for MLP layers. Analytically, the additional parameters can be calculated by $L\times d\times 128+L\times 128\times(\frac{d}{H})+L\times d\times N$. Let’s use LLaMA-2 7B as an example, $L=32$, $d=4096$, $N=8$, $H=32$, the additional parameters are $1\times 32\times 4096\times 128+32\times 128\times(128)+32\times 4096\times 8=0.0184$B. This is equivalent to $0.27\%$ of the total parameters of the LLaMA-2 7B model, and thus, the additional parameter is not significant compared to the original number of parameters.
@@
Rotary Position Embedding (RoPE) (su2024roformer) is a popular positional encoding method, and it is regularly used in LLMs like LLaMA (touvron2023llama2). RoPE divides the $\frac{d}{H}$ dimensional space into $\frac{d}{2H}$ sub-spaces, and they are applied on query and key. This means that if we want to perform head dimension pruning for query and key matrices, we need to follow the sub-spaces resulting from RoPE and make these two sub-spaces share the same pruning mask $\mathbf{s}_{0}$, and the final pruning mask for query and key is $\mathbf{s}_{0}^{\prime}=[{\mathbf{s}_{0}}_{[1:\frac{d}{2H}]},{\mathbf{s}_{0}}_{[1:\frac{d}{2H}]}]$, and clearly the size of ${\mathbf{s}_{0}}_{[1:\frac{d}{2H}]}$ is $\frac{d}{2H}$. In short, we simply select the first half of elements from $\mathbf{s}_{0}$ and repeat it twice to make the final pruning decision. We also found that applying dynamic pruning for query and key matrices along the head dimension is difficult and unreasonable since different tokens may have different positions after pruning. It becomes a problem when calculating the inner product between the query and key matrices given different tokens.
By applying head dimension pruning, our method also does not need to be modified when facing different attention mechanisms like GQA (Grouped-Query Attention) (ainslie2023gqa) and MQA (Multi-Query Attention) (shazeer2019fast).
@@
70% SliceGPT (ashkboos2023slicegpt) 60.27 36.18 69.42 64.09 49.74 55.94
LLM Surgeon (van2023llm) 69.74 40.27 76.50 68.67 71.52 65.34
DISP-LLM (gaodisp) 63.80 39.42 74.43 66.85 70.86 63.07
DISP-LLM Alpaca (gaodisp) 68.98 44.28 77.31 67.32 68.98 65.59
MoDeGPT-Alpaca (lin2024modegpt) 70.24 41.47 77.15 71.27 71.84 66.39
@@
60% SliceGPT (ashkboos2023slicegpt) 48.99 32.51 63.17 56.75 39.85 48.25
LLM Surgeon (van2023llm) 63.80 37.12 73.16 65.75 65.64 60.94
DISP-LLM (gaodisp) 62.67 35.63 73.39 62.67 65.86 60.04
DISP-LLM Alpaca (gaodisp) 66.79 42.75 75.30 64.25 67.52 63.32
MoDeGPT-Alpaca (lin2024modegpt) 63.72 38.82 71.87 66.30 62.10 60.56
@@
50% LLM Surgeon (van2023llm) 56.19 37.12 68.87 63.22 56.19 56.32
DISP-LLM (gaodisp) 58.27 36.87 68.67 59.27 57.18 54.50
DISP-LLM Alpaca (gaodisp) 55.72 37.54 72.20 59.59 62.39 57.49
@@ A.3 Details of Gumbel-Softmax and Gumbel-Sigmoid
The Gumbel-Softmax function (jang2016categorical) allows for differentiable sampling from a categorical distribution. Given logits $\mathbf{x}$, the Gumbel-Softmax sample $\mathbf{y}$ is computed as:
@@
where each element of $\mathbf{g}$ is drawn from $\text{Gumbel}(0,1)$, and $\tau$ is the temperature parameter that controls the smoothness of the distribution. Combining Gumbel-Softmax with the Straight-Through gradient Estimator (bengio2013estimating), we have the following equation:
$\text{ST-GSmax}(\mathbf{x})=\text{one-hot}\left(\arg\max_{i\in D}\left[\frac{x_{i}+g_{i}}{\tau}\right]\right)$ (14)
where $D=\{1,2,\cdots,N\}$, $N$ again is the number of experts in our setting, and one-hot will assign $1$ corresponding to the position of the maximum value in $\frac{\mathbf{x}+\mathbf{g}}{\tau}$ and assign 0 to other positions.
The Gumbel-Sigmoid function is a special case of the Gumbel-Softmax function, designed for binary distributions. Given logits $\mathbf{x}$, the Gumbel-Sigmoid sample $\mathbf{y}$ is computed as:
$\mathbf{y}=\text{sigmoid}\left(\frac{\mathbf{x}+\mathbf{g}}{\tau}\right),$
@@
$\text{ST-GSig}(\mathbf{x})=\text{round}(\text{sigmoid}\left(\frac{\mathbf{x}+\mathbf{g}+b}{\tau}\right)),$ (15)
where $b$ is a constant bias in our implementation and it ensures that all experts start from the whole model, $\text{round}(\cdot)$ will round the input values to the nearest integer, and in our case, it rounds inputs to 0 or $1$. For all experiments, we set $b=3.0$ in Eq. 15, and we set $\tau=0.4$ for Eq. 14 and Eq. 15.
@@
DISP-LLM (gaodisp) 68.18 44.11 74.86 67.09 62.93 63.43
75% SliceGPT (ashkboos2023slicegpt) 53.70 31.66 69.21 65.35 52.40 54.46
@@
DISP-LLM (gaodisp) 65.93 43.34 74.27 65.11 59.95 61.72
70% SliceGPT (ashkboos2023slicegpt) 53.03 30.29 65.94 63.14 47.56 51.99
@@
Appendix B More Details of the Loss Design
B.1 Implementation of the Self-Knowledge Distillation
During the ToMoE learning process, we freeze the parameters of the original model. This approach offers the additional benefit of enabling self-knowledge distillation without the need to load an extra model. In Lst. 1, we present the pseudo-code for the self-knowledge distillation process. In summary, we first disable the trainable modules associated with ToMoE and compute the output logits from the original model. Next, we re-enable the trainable modules for ToMoE and perform a regular forward pass. The logits from the original model are then used to guide the learning of ToMoE.
@@ B.1 Implementation of the Self-Knowledge Distillation
Table 12: Zero-shot task performance of compressed LLaMA-2 7B with more settings.
Active Parameters Method ARC-e ARC-c PIQA WinoG. HellaS. Average
acc-norm acc-norm acc-norm acc acc-norm
100% Dense 74.58 46.25 79.11 69.06 75.99 69.00
70% ShortGPT (men2024shortgpt) 48.65 32.85 64.31 64.33 56.13 53.25
SliceGPT (ashkboos2023slicegpt) 58.88 33.36 68.55 58.01 49.86 53.73
LLM Surgeon (van2023llm) 63.09 36.69 73.56 61.09 60.72 59.03
DISP-LLM  (gaodisp) 59.81 33.19 71.82 62.27 63.43 58.10
DISP-LLM Alpaca (gaodisp) 60.10 37.03 73.72 63.93 62.87 59.53
ModeGPT (lin2024modegpt) 63.26 38.73 70.40 67.32 63.26 60.78
ModeGPT-Alpaca (lin2024modegpt) 65.49 39.16 73.34 66.22 65.90 62.02
60% ShortGPT (men2024shortgpt) 41.16 29.94 60.12 60.46 43.67 47.07
SliceGPT (ashkboos2023slicegpt) 36.49 24.57 54.90 53.43 34.80 40.84
LLM Surgeon (van2023llm) 52.31 30.29 69.26 54.38 48.04 50.86
ModeGPT (lin2024modegpt) 49.45 30.03 64.96 61.96 53.01 51.88
ModeGPT-Alpaca (lin2024modegpt) 59.76 34.73 70.35 64.40 58.63 57.58
ToMoE (Ours) 63.64 38.74 72.85 62.51 65.84 60.72
50% LLM Surgeon  (van2023llm) 44.91 26.28 64.36 52.57 40.29 45.68
DISP-LLM  (gaodisp) 43.06 25.85 63.93 54.54 63.43 46.72
DISP-LLM Alpaca (gaodisp) 51.14 30.20 68.34 56.20 49.35 51.05
ToMoE (Ours) 56.65 33.87 71.00 58.56 60.26 56.07